# Drift time calibration of the ultra-low material budget GEM-based TPC for MIXE

### v2

@@ 1 Introduction
Muon-Induced X-ray Emission (MIXE) [1, 2, 3, 4] is a powerful technique that leverages exotic atom formation by irradiating target samples with a well-defined beam of negative muons. As muons stop within the material, they are captured by target atoms, forming excited atomic states and cascading to the ground state on a picosecond time scale [5, 6]. During this process, X-rays are emitted at characteristic energies specific to the atomic species, enabling precise identification of both elemental and isotopic compositions. In addition to atomic cascades, muons can also undergo nuclear capture, a process that becomes more probable in heavier elements [7]. During nuclear capture, the muon induces the conversion of a proton into a neutron, resulting in a transmuted nucleus. These nuclei de-excite by emitting a combination of nucleons and gamma rays. The combination of muonic X-rays from atomic cascades and gamma rays from nuclear interactions provides multiple pathways for elemental identification. High-purity germanium (HPGe) detectors are used to measure these emissions with excellent energy resolution and efficiency over a broad spectrum, ensuring accurate and detailed analysis. One significant advantage of MIXE over non-destructive techniques like X-ray fluorescence (XRF) [8] is its depth sensitivity. By adjusting the momentum of implanted muons, their stopping range is well defined spanning micrometers to centimeters, enabling advanced non-destructive, depth-resolved elemental analysis. Thus, MIXE is highly suitable for assessing valuable, rare or fragile samples, making it ideal for studying cultural heritage artifacts [9] or for operando investigations such as battery charging kinetics [10]. To further enhance the method, we aim to correlate the precise muon stopping location with the corresponding X-ray emissions, thereby enabling a novel, element-sensitive tomographic technique.
@@ 2 Drift time calibration
where the model assumes that the Gaussian peaks are symmetrically distributed around the central time4 $C$, with offset timing $D$. The parameters $A_{1},~{}A_{2}$ and $A_{3}$ represent the amplitudes and $W_{1}$, $W_{2}$ and $W_{3}$ represent the widths of the respective distributions. Since the separation between the fibers directly corresponds to the offset timing, it is then simply given by $v_{d}=4\,\mathrm{mm}/{\left|D\right|}$.

### v3

@@ 1 Introduction
Muon-Induced X-ray Emission (MIXE) [1, 2, 3, 4] is a powerful technique that leverages exotic atom formation by irradiating target samples with a well-defined beam of negative muons. As muons stop within the material, they are captured by target atoms, forming excited atomic states and cascading to the ground state on a picosecond time scale [5, 6]. During this process, X-rays are emitted at characteristic energies specific to the atomic species, enabling precise identification of both elemental and isotopic compositions. In addition to atomic cascades, muons can also undergo nuclear capture, a process that becomes more probable in heavier elements [7]. During nuclear capture, the muon induces the conversion of a proton into a neutron, resulting in a transmuted nucleus. These nuclei de-excite by emitting a combination of nucleons and gamma rays. The combination of muonic X-rays from atomic cascades and gamma rays from nuclear interactions provides multiple pathways for elemental identification. High-purity germanium (HPGe) detectors are used to measure these emissions with excellent energy resolution and efficiency over a broad spectrum, ensuring accurate and detailed analysis. One significant advantage of MIXE over other non-destructive techniques, like X-ray fluorescence (XRF) [8], is its depth sensitivity. By adjusting the momentum of implanted muons, their stopping range is well defined, spanning micrometers to centimeters, enabling advanced non-destructive, depth-resolved elemental analysis. Thus, MIXE is highly suitable for assessing valuable, rare, or fragile samples, making it ideal for studying cultural heritage artifacts [9] or for operando investigations such as battery charging kinetics [10]. To further enhance the method, we aim to correlate the precise muon stopping location with the corresponding X-ray emissions, thereby enabling a novel, element-sensitive tomographic technique.
@@ 2 Drift time calibration
where the model assumes that the Gaussian peaks are symmetrically distributed around the central time $C$, with offset timing $D$. The parameters $A_{1},~{}A_{2}$ and $A_{3}$ represent the amplitudes and $W_{1}$, $W_{2}$ and $W_{3}$ represent the widths of the respective distributions. Since the separation between the fibers directly corresponds to the offset timing, it is then simply given by $v_{d}=4\,\mathrm{mm}/{\left|D\right|}$.